### v2

# MCGA: A Multi-task Classical Chinese Literary Genre Audio Corpus

Abstract
With the rapid advancement of Multimodal Large Language Models (MLLMs), their potential has gained significant attention in Chinese Classical Studies (CCS). While existing research primarily focuses on text and visual modalities, the audio corpus within this domain remains largely underexplored. To bridge this gap, we introduce the Multi-task Classical Chinese Literary Genre Audio Corpus (MCGA), a 119-hour corpus comprising 22,000 audio samples. It encompasses a diverse range of literary genres across six tasks: Automatic Speech Recognition (ASR), Speech-to-Text Translation (S2TT), Speech Emotion Captioning (SEC), Spoken Question Answering (SQA), Speech Understanding (SU), and Speech Reasoning (SR). Through the evaluation of ten MLLMs, our experimental results demonstrate that current MLLMs still face substantial challenges on the MCGA test set. Furthermore, we introduce a domain-specific metric for SEC and a metric to measure the consistency between speech and text capabilities. We release MCGA to the public to facilitate the development of more robust MLLMs. [^1] MCGA Corpus: https://github.com/yxduir/MCGA
MCGA: A Multi-task Classical Chinese Literary Genre Audio Corpus
Yexing Du1,2   Kaiyuan Liu1,2   Bihe Zhang3   Youcheng Pan2   Bo Yang2   Liangyu Huo4 Xiyuan Zhang4   Jian Xie4   Daojing He1   Yang Xiang2   Ming Liu1,2   Bing Qin1,2 1Harbin Institute of Technology  2Pengcheng Laboratory 3South China University of Technology  4Du xiaoman {yxdu, mliu}@ir.hit.edu.cn, {panych, xiangy}@pcl.ac.cn
1 Introduction
The development of Multimodal Large Language Models (MLLMs) has significantly advanced Chinese Classical Studies (CCS). These models support multimodal inputs, providing powerful capabilities for interpreting ancient texts, which in turn enhances cultural preservation Zhang et al. (2025). However, while most existing research focuses on textual Cao et al. (2024) or visual Liu et al. (2025b) modalities, the auditory dimension of CCS remains largely unexplored. This gap stems from a lack of high-quality, domain-specific audio corpora, thereby constraining the potential for an omni-modal understanding of CCS.
Figure 1: Timeline of the Golden Age for Classical Chinese Literary Genres: Fu (Rhapsody), Shi (Poetry), Wen (Prose), Ci (Lyric), and Qu (Song).
Figure 2: Examples from the MCGA Corpus. The corpus covers six core speech tasks (ASR, S2TT, SEC, SQA, SU, SR). Leveraging its parallel speech-text data, it also supports four text tasks: Machine Translation (MT), Question Answering (QA), Language Understanding (LU), and Language Reasoning (LR).
To bridge this critical gap, we introduce the Multi-task Classical Chinese Literary Genre Audio Corpus (MCGA), a comprehensive resource designed to catalyze audio-centric research in CCS. As illustrated in Figure 1, MCGA encompasses five primary literary genres: Fu, Shi, Wen, Ci, and Qu. The corpus consists of 22,000 audio samples, totaling 119 hours of recorded content. To ensure cultural and linguistic authenticity, the data were recorded by native speakers in standard Mandarin Chinese. Crucially, all audio samples include explicit copyright transfers, thereby resolving long-standing Intellectual Property Rights (IPR) challenges in open-source CCS audio datasets.
The MCGA corpus offers two primary advantages: (1) Task Diversity: As illustrated in Figure 2, the corpus supports 6 diverse speech-centric tasks, including Automatic Speech Recognition (ASR), Speech-to-Text Translation (S2TT), Speech Emotion Captioning (SEC), Spoken Question Answering (SQA), Speech Understanding (SU), and Speech Reasoning (SR), alongside four integrated text tasks. (2) Literary Genre Diversity: It encompasses 5 major literary genres spanning 11 historical periods, forming a total of 37 distinct period-genre categories and covering a comprehensive collection of 4,497 literary works.
We evaluated 10 representative MLLMs, including 2 closed-source and 8 open-source models. Experimental results indicate that current MLLMs still have significant room for improvement in the CCS field. Notably, even the top-performing model, Qwen3-Omni Xu et al. (2025b), scored below 60 on complex tasks such as SEC. Besides, we introduce a novel evaluation metric tailored for literary SEC, along with a Cross-Modal Consistency (CMC) metric to quantify the alignment between a model’s auditory and textual reasoning. Furthermore, the substantial performance gains achieved through training underscore the superior quality of the MCGA corpus.
Dataset (Text / Image / Audio) Modality Domain Scale License Copyright
ACLUE Zhang and Li (2023) CCS 4,967 CC BY-NC-4.0
CCLUE Wang et al. (2023) CCS 36,319 Apache-2.0
WYWEB Zhou et al. (2023) CCS 69,700 -
WenMind Cao et al. (2024) CCS 4,875 CC BY-NC-SA-4.0
TianWen Pei et al. (2025) CCS 4,000 MIT
CII-Bench Zhang et al. (2025) / General 698 Apache-2.0
FoodieQA Li et al. (2024) / Food 389 CC BY-NC-ND-4.0
Oracle-Bench Qiao et al. (2025) / CCS 2,834 -
Paint4Poem Li et al. (2021) / CCS 93,153 Github
MCS-Bench Liu et al. (2025b) / CCS 6,500 CC BY-NC-SA-4.0
MCGA (ours) / CCS 22,000 CC BY-NC-SA-4.0
Table 1: Comparison of MCGA with Existing Chinese Cultural Datasets. MCGA is the first large-scale, fully copyrighted classical Chinese literary audio corpus for MLLMs (119 hours). All recordings are sourced directly from original creators with full copyright transfer, highlighting our commitment to Intellectual Property Rights (IPR) protection in Chinese Classical Studies (CCS) research.
Our primary contributions are as follows:
  • MCGA Corpus: We present MCGA, the first large-scale (119 hours), open-source, and fully copyrighted audio corpus dedicated to classical Chinese literature studies. This resource effectively bridges the gap in high-quality audio datasets for this domain.
  • Evaluation Framework: We establish a comprehensive evaluation framework centered on MCGA, comprising 6 multifaceted tasks: ASR, S2TT, SEC, SQA, SU, and SR. This enables a rigorous investigation into the capabilities of MLLMs.
  • Evaluation Metrics: We introduce two novel evaluation metrics: a domain-specific metric tailored for the literary SEC task, and a CMC metric designed to assess the alignment and quantify the gap between auditory and textual modalities.
  • Empirical Analysis: We evaluate 10 MLLMs to identify performance bottlenecks in the classical Chinese literature domain. Besides, we demonstrate MCGA’s high utility as a training resource, where fine-tuning yields substantial performance breakthroughs.
2 Related Works
2.1 Chinese Cultural Datasets
The landscape of Chinese cultural evaluation spans many domains. ACLUE Zhang and Li (2023) and WYWEB Zhou et al. (2023) establish large-scale benchmarks for Classical Chinese and ancient literature, focusing on linguistic understanding. Complementarily, CCLUE Wang et al. (2023) rethinks cultural evaluation across broader contexts. In the multimodal sphere, FoodieQA Li et al. (2024) and CII-Bench Zhang et al. (2025) probe culinary arts and figurative reasoning, respectively, highlighting a shift toward assessing complex cultural heritage and everyday traditions.
2.2 Chinese Classical Studies Datasets
Recent benchmarks deepen the evaluation of Chinese classical heritage through diverse methodologies. WenMind Cao et al. (2024) assesses deep cultural cognition and mentalities, while TianWen Pei et al. (2025) provides specialized assessment for traditional scriptures and historical knowledge. Advancing into multimodality, Oracle-Bench Qiao et al. (2025) evaluates ancient script deciphering, whereas Paint4Poem Li et al. (2021) bridges classical poetry with visual synthesis. MCS-Bench Liu et al. (2025b) offers a framework for multimodal classical studies. However, few of these benchmarks or datasets contain the parallel speech of the classical Chinese literature.
3 MCGA Corpus
3.1 Overview
We introduce MCGA, a comprehensive corpus designed to promote audio-centric research in CCS. This section briefly outlines the construction, the human recording process, the subsequent quality control and the statistics of MCGA.
3.2 Data Construction
Data Collection and Preprocessing.
Classical Chinese literature and corresponding Pinyin were sourced from the web. All works are in the public domain (created over 150 years ago). Following rigorous cleaning, texts were segmented by sentence boundaries and character counts to limit recording lengths to under 30 seconds.
Text Data Construction.
Subsequently, we leverage DeepSeek-V3.2 Guo et al. (2025) to generate question-answer pairs for the text of each clip, with access to the full literary context. This process covers a variety of speech-related tasks, including S2TT, SEC, SQA, SU, and SR.
Text Data Verification.
The generated question–answer pairs are subjected to trio validation using DeepSeek-V3.2, GPT-5-mini OpenAI (2025), and Gemini-3-Flash Team and DeepMind (2025), through which pairs that fail to pass the verification are filtered out. The test and validation sets underwent human verification to ensure data quality.
Figure 3: MCGA Corpus Construction. Initially comprising only metadata such as titles, authors, and texts, the MCGA corpus is expanded through human recording, LLM generation, and rigorous verification. Then, it supports six speech tasks: ASR, S2TT, SEC, SQA, SU, and SR. We provide a detailed example of the SEC task in Figure 4.
3.3 Human Recording
Volunteer Demographics.
We recruited 28 native speakers (13 males and 15 females, aged 18–40) to record the texts via a dedicated private website. All participants have good educational backgrounds, half of whom are Chinese majors.
Recording Protocol.
We explicitly stated the recording guidelines to the volunteers, as follows:
Volunteer Guidelines use a tone that matches the emotion of the text. can access the pinyin for all Chinese characters. ensure the recording environment is quiet. each clip must be read in 30 seconds. each clip is read by at least 1 male and 1 female. clips from the same work are sent to the same person.
3.4 Audio Quality Check
MLLM Verification.
We employed a dual-stage speech recognition verification process using Qwen and Whisper Radford et al. (2023) models to identify samples with significant errors, which were subsequently re-recorded by the volunteers.
Human Verification.
We recruited 6 data quality inspection volunteers to verify the validation and test sets. The inspectors were instructed to score the samples. Low quality samples (pronunciation error or presence of background noise) were removed from the sets.
Both recording volunteers and quality inspection volunteers signed labor agreements and were compensated with reasonable remuneration.
Figure 4: Case for SEC Task.
Figure 5: Corpus Statistics. It comprises 22,000 filtered human-recorded speech samples (totaling 119 hours) and supports 6 downstream tasks. Sample counts for S2TT, SEC, SU, and SR are lower due to the removal of invalid QA pairs. (NSD: the Northern and Southern Dynasties; FD: the Five Dynasties)
3.5 Case Study for SEC
SEC Task.
To capture the emotional and artistic nuances of classical literature, we present a case study in Figure 4. For SEC, the MLLM must sequentially generate three components:
  1. 1.
    Persona Profiles: Analysis of the speaker, including age and gender.
  2. 2.
    Overall Sentiment Analysis: A summary of the general emotional tone and attitude.
  3. 3.
    Sentence Transcription and Emotion: A sentence-by-sentence decomposition in the format: “Transcription | Emotion.”
3.6 Dataset Statistics
Figure 5 shows the statistics of MCGA. It spans 5 genres across 11 historical periods, resulting in 37 unique period–genre categories.
Tang Shi has the most samples, followed by Song Ci. This is because the Tang and Song dynasties were the two peak periods of classical Chinese literature. Shi was the most popular genre in the Tang dynasty, while Ci was in the Song dynasty.
The corpus comprises 22,000 filtered human-recorded audio samples (totaling 119 hours) and supports 6 downstream tasks: ASR, S2TT, SEC, SQA, SU, and SR. The longest audio sample is 30 seconds, the shortest is 3.5 seconds, and the average duration is 19.5 seconds.
It should be noted that sample counts for S2TT, SEC, SU, and SR are lower due to the removal of invalid QA pairs. Also, the validation or test sets for the six tasks are not parallel.
Task Metric Details
ASR CER $\downarrow$ Text normalization
Following Morris et al. (2004)
S2TT LLM-B $\uparrow$ LLM Evaluation
Following Chen et al. (2025)
SEC LLM-C $\uparrow$ LLM Evaluation
Proposed in this work
SQA F1 $\uparrow$ Open-ended
Factuality Evaluation
SU Accuracy $\uparrow$ Multiple-choice questions
Options derived from the speech
SR Accuracy $\uparrow$ Multiple-choice questions
External knowledge reasoning
Table 2: Metric Details.
4 Experiments
4.1 Experiment Setting
Baseline MLLMs.
We evaluate 2 closed-source MLLMs (GPT-4o-mini-Audio OpenAI (2023) and Gemini-3-Flash Team and DeepMind (2025)) and 8 open-source MLLMs: the Qwen series Chu et al. (2024); Xu et al. (2025a, b), the Voxtral series Liu et al. (2025a), Phi-4-Multimodal-Instruct Abouelenin et al. (2025), MiDashengLM Dinkel et al. (2025), and Step-Audio-2-mini Wu et al. (2025).
Training Details.
We fine-tuned Qwen2.5-Omni-7B using the ms-swift framework [^2] https://github.com/modelscope/ms-swift with LoRA ($r=8,\alpha=32$Hu et al. (2021). The model was trained for 3 epochs on 4 A100 GPUs using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, a per-device batch size of 8, and a gradient accumulation of 4.
Evaluation Metrics.
As shown in Table 4, we evaluate MLLMs across six tasks. All open-source models are deployed using the vLLM framework [^3] https://github.com/vllm-project/vllm Kwon et al. (2023), with inference performed via API requests at a temperature of 0. To provide a more intuitive performance metric, we normalize the S2TT and SEC results to a 100-point scale. Specifically, the ASR task is evaluated using the Character Error Rate (CER) [^4] https://github.com/jitsi/jiwer, while the S2TT and SEC tasks are scored by the deepseek-chat API Guo et al. (2025). For SQA, we report the F1 score, and for SU and SR, we report Accuracy.
Model ASR S2TT SEC SQA SU SR
CER $\downarrow$ LLM-B $\uparrow$ LLM-C $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ Acc $\uparrow$
Closed-source MLLMs
GPT-4o-mini-Audio OpenAI (2023) 20.6 43.5 5.7 30.6 74.8 70.2
Gemini-3-Flash Team and DeepMind (2025) 6.1 74.0 54.0 48.7 86.6 83.7
Open-source MLLMs
Phi-4-Multimodal-Instruct Abouelenin et al. (2025) 59.6 27.5 12.7 24.5 50.6 54.4
Voxtral-Mini Liu et al. (2025a) 30.0 25.5 15.0 12.5 58.9 62.8
Voxtral-Small Liu et al. (2025a) 30.0 34.1 16.4 27.9 72.6 71.9
MiDashengLM Dinkel et al. (2025) 11.7 42.9 24.7 22.5 72.2 75.6
Step-Audio-2-mini Wu et al. (2025) 9.9 41.9 36.8 45.2 80.5 80.4
Qwen2-Audio-7B-Instruct Chu et al. (2024) 18.7 30.4 26.1 24.8 72.1 64.7
Qwen2.5-Omni-7B Xu et al. (2025a) 10.1 49.7 37.0 43.5 81.3 79.3
Qwen3-Omni-30B-A3B-Instruct Xu et al. (2025b) 4.4 67.6 58.4 51.5 86.9 82.9
Table 3: Performance Comparison of Different MLLMs on the MCGA Test Set. Detailed results for LLM-B and LLM-C are shown in Table 6 and Table 7.
4.2 Main Results
We present a comprehensive evaluation of ten MLLMs across six audio tasks. By analyzing the interplay between model performance and task difficulty, we derive the following key observations:
Closed-source vs. Open-source Models.
In Table 3, Qwen3-Omni demonstrates superior performance on the MCGA test set for Chinese understanding and generation tasks, specifically in ASR (4.4 CER $\downarrow$), SEC (58.4 LLM-C $\uparrow$), SQA (51.5 F1 $\uparrow$), and SU (86.9 Acc $\uparrow$). Conversely, closed-source models such as Gemini-3-Flash maintain a competitive edge in English generation and Chinese reasoning tasks, leading in metrics such as S2TT (74.0 LLM-B $\uparrow$) and SR (83.7 Acc $\uparrow$). Overall, the open-source models have achieved a competitive level of performance compared with the closed-source ones.
Comparison across Different Tasks.
As shown in Figure 6, existing MLLMs demonstrate their strongest performance in CCS ASR tasks. This is followed by SU and SR in multiple-choice formats, where models achieve relatively robust results. Regarding the S2TT task, overall performance is acceptable but remains to be enhanced. In contrast, performance on SEC is notably poor, indicating a critical need for enhanced affective computing capabilities. Finally, for SQA, F1 scores remain low, suggesting that "hallucination" issues Du et al. (2025) have yet to be effectively resolved.
Figure 6: Comparison across Different Tasks. Existing MLLMs exhibit robust performance in ASR, SU, and SR tasks, but they still encounter challenges regarding the beauty of translation in S2TT, affective modeling in SEC, and hallucination issues in open-ended SQA. CER refers to $(1-\text{CER}\%)$.
4.3 Further Analysis
4.3.1 Analysis of ASR Task
Performance Disparity Across Genres.
Table 4 shows the MLLMs’ performance which varies significantly by genre. Qwen3-Omni achieves state-of-the-art results on MCGA, maintaining the lowest CER across all categories, particularly in Ci (2.8). A consistent trend across various models is that Ci achieves lower CER, while Fu consistently poses the greatest difficulty. This difficulty stems from Fu’s ornate rhetoric, frequent classical allusions, and high density of modal particles.
Models Shi Ci Qu Fu Wen
GPT-4o-mini-Audio 22.6 20.3 19.0 22.8 18.5
Gemini-3-Flash 6.1 6.8 7.7 8.4 6.1
Phi-4-Multimodal-Instruct 61.0 63.8 63.0 61.1 51.1
Voxtral-Mini 29.2 27.4 29.2 34.5 28.9
Voxtral-Small 30.6 26.4 30.1 33.9 28.4
MiDashengLM 12.7 10.1 9.4 15.7 10.0
Step-Audio-2-mini 9.0 7.0 7.5 14.6 10.3
Qwen2-Audio-7B-Instruct 18.2 15.8 15.9 23.3 19.0
Qwen3-Omni-30B-A3B-Instruct 3.8 2.8 4.1 6.2 4.3
Qwen2.5-Omni-7B 9.9 7.5 8.9 14.8 8.8
Qwen-Omni-MCGA 2.8 3.1 7.8 5.3 4.1
Table 4: CER Scores Across Different Genres. The test set contains 1,000 samples (200 per genre). Underline indicates the best-performing genre for each individual model. Qwen-Omni-MCGA is a LoRA-based adaptation of Qwen2.5-Omni-7B. It achieves state-of-the-art results on all genres except for Qu.
Figure 7: CER Across Dynasties and Genres.
Audio Quality.
Table 5 reveals a 0.1 CER gap (Qwen3-Omni) between human-verified valid/test sets and the train set, confirming high data consistency. Residual errors primarily stem from uncommon characters and phonetic loanwords (tongjiazi) in Classical Chinese. Figure 7 shows the CER distribution across dynasties and genres.
Models Train Valid Test Avg.
MiDashengLM 11.7 11.2 11.7 11.7
Step-Audio-2-mini 10.4 10.1 9.9 10.4
Qwen2.5-Omni-7B 9.8 9.5 10.1 9.8
Qwen3-Omni-30B-A3B-Instruct 4.5 4.4 4.4 4.5
Table 5: CER Scores for Quality Check. The train, valid, and test sets show high data consistency.
4.3.2 Analysis of S2TT Task
Beauty Evaluation of Translation.
As illustrated in Table 6, we evaluate the translation quality across four dimensions: Beauty of Form (LLM-BF), Beauty of Meaning (LLM-BM), Beauty of Sound (LLM-BS), and their average score (LLM-B). The closed-source model Gemini-3-Flash achieves the highest performance across all metrics, reaching a peak average score of 74.0 (LLM-B $\uparrow$).
S2TT Quality.
Additionally, we provide high-quality ground-truth translation candidates. The LLM-B score of 79.2 (4.0) is constrained by the 1–5 evaluation scale, as the DeepSeek API evaluation model typically assigns moderate scores and rarely grants a perfect score of 5. To provide a more intuitive performance metric, we normalize these raw API scores to a 100-point scale.
Models LLM-BF LLM-BM LLM-BS LLM-B
Ground Truth 79.8 80.3 77.4 79.2
GPT-4o-mini-Audio 42.3 42.8 45.3 43.5
Gemini-3-Flash 72.4 74.5 75.2 74.0
Phi-4-Multimodal-Instruct 27.0 27.3 28.1 27.5
Voxtral-Mini 25.6 25.6 25.1 25.5
Voxtral-Small 34.5 34.2 33.6 34.1
MiDashengLM 43.8 44.3 40.6 42.9
Step-Audio-2-mini 42.6 43.3 39.6 41.9
Qwen2-Audio-7B-Instruct 29.8 30.6 30.9 30.4
Qwen2.5-Omni-7B 51.3 51.7 46.1 49.7
Qwen3-Omni-30B-A3B-Instruct 68.1 69.1 65.5 67.6
Table 6: Beauty Evaluation of Translation. Following Chen et al. (2025), we employ Beauty of Form (BF), Beauty of Meaning (BM) and Beauty of Sound (BS) as evaluation metrics. LLM-B denotes the mean of these three evaluation metrics.
4.3.3 Analysis of SEC Task
SEC Evaluation.
We design an LLM-based penalty evaluation mechanism based on reference answers. The mechanism consists of the following three metrics:
  • Persona Recognition (SEC-P, 0–2): Measures the capability to extract identity features such as age and gender. Starting from an initial score of $2$, a penalty of $1$ point is deducted for each attribute error.
  • Global Emotional Tone (SEC-G, 0–3): Evaluates the overall emotional atmosphere based on the richness and accuracy of the descriptions. A score of 0 is assigned if the emotional category or context is misidentified.
  • Sentence-level Emotion Tracking (SEC-S, 0–5): Evaluates sentence-by-sentence transcription and analysis. 1 point is deducted for each error in emotional portrayal. If the transcription is entirely unrelated (hallucination), a score of 0 is recorded.
Open-source vs. Closed-source MLLMs.
As shown in Table 7, Qwen3-Omni outperforms other models across all SEC metrics. This superior performance is attributed to its deep understanding of Chinese cultural nuances and its robust transcription capabilities. It is followed by Gemini-3-Flash, which maintains competitive results.
In contrast, GPT-4o-mini-Audio exhibits poor performance. This is primarily because its stringent safety protocols frequently trigger refusals when tasked with persona-based or emotional analysis.
Models SEC-P SEC-G SEC-S LLM-C
Ground Truth 20.0 30.0 50.0 100.0
GPT-4o-mini-Audio 1.4 3.2 1.2 5.7
Gemini-3-Flash 13.4 16.9 23.6 54.0
Phi-4-Multimodal-Instruct 4.3 7.9 0.5 12.7
Voxtral-Mini 5.5 8.9 0.6 15.0
Voxtral-Small 2.2 11.9 2.3 16.4
MiDashengLM 13.6 8.0 3.2 24.7
Step-Audio-2-mini 16.2 11.8 8.8 36.8
Qwen2-Audio-7B-Instruct 12.9 10.4 2.7 26.1
Qwen2.5-Omni-7B 14.3 13.9 8.8 37.0
Qwen3-Omni-30B-A3B-Instruct 16.0 18.6 23.8 58.4
Table 7: LLM-based Evaluation for SEC. (1) SEC-P ($0\text{--}2$) for persona identification; (2) SEC-G ($0\text{--}3$) for global emotional tone analysis; (3) SEC-S ($0\text{--}5$) for sentence-level emotion; (4) LLM-C is the sum of scores.
4.3.4 Analysis of SQA, SU, and SR Tasks
Open-ended vs. Multiple-choice QA.
As shown in Table 8, a substantial performance gap exists between multiple-choice and open-ended formats. MLLMs struggle significantly more with open-ended questions, such as identifying authors or titles, compared to complex reasoning tasks that provide candidate options. For instance, Gemini-3-Flash scored 86.6 in SU and 83.7 in SR but drops to 48.7 in SQA. This gap indicates that MLLMs suffer from severe hallucinations in open-ended factual QA, despite their strong reasoning.
Cross-modal Consistency.
To evaluate how reliably MLLMs maintain consistency across different input modalities, we define the Cross-modal Consistency (CMC) metric as:
$\mathrm{CMC}=\frac{1}{3}\left(\frac{\mathrm{SQA}}{\mathrm{QA}}+\frac{\mathrm{SU}}{\mathrm{LU}}+\frac{\mathrm{SR}}{\mathrm{LR}}\right)\times 100$ (1)
As shown in Table 8, SQA, SU, and SR represent performance on speech-based tasks, while the denominators QA, LU (Language Understanding), and LR (Language Reasoning) serve as the text-only upper-bound references. Step-Audio-2-mini achieved the highest CMC score among all evaluated MLLMs.
Models SQA SU SR QA LU LR CMC
Gemini-3-Flash 48.7 86.6 83.7 66.0 94.6 91.5 85.6
Phi-4-Multimodal-Instruct 24.5 50.6 54.4 25.2 69.1 60.4 86.8
Voxtral-Mini 12.5 58.9 62.8 13.7 76.8 69.5 86.1
Voxtral-Small 27.9 72.6 71.9 40.0 88.8 83.1 79.3
MiDashengLM 22.5 72.2 75.6 45.1 89.1 84.4 73.5
Step-Audio-2-mini 45.2 80.5 80.4 51.3 89.6 85.4 90.7
Qwen2-Audio-7B-Instruct 24.8 72.1 64.7 36.0 79.9 70.7 83.5
Qwen2.5-Omni-7B 43.5 81.3 79.3 53.1 91.6 85.0 88.0
Qwen3-Omni-30B-A3B-Instruct 51.5 86.9 82.9 60.0 93.6 91.0 89.9
Table 8: Cross-modal Consistency. CMC quantifies the performance gap between audio and textual modalities.
5 Conclusion
This paper presents MCGA, the first large-scale, fully copyrighted audio corpus dedicated to CCS, encompassing six tasks. Besides, we propose a domain-specific evaluation metric for SEC alongside a metric for cross-modal consistency. Our systematic evaluation of ten MLLMs reveals that the Qwen series leads in performance, exhibiting superior proficiency in comprehending classical Chinese literature.
6 Limitations
Although MCGA incorporates audio-text multimodal data across six distinct tasks, several limitations persist. First, copyright constraints preclude the inclusion of real-world image samples that are precisely aligned with both textual and auditory modalities. Second, the Qu genre emerged significantly later than Shi, Ci, Wen, and Fu. Due to the relatively short-lived nature of the Yuan Dynasty, the volume of extant works is considerably limited, leading to a lower representation of Qu within the corpus.
7 Ethical Considerations
We emphasize that ethical standards are of paramount importance in research involving human audio data. All audio data used in this study were recorded by human volunteers who contacted the authors directly. The volunteers were fairly compensated for their contributions and have signed a Voice Authorization License Agreement, explicitly granting permission for their recorded speech to be used for research purposes. Data handling and usage strictly comply with all applicable privacy and data protection regulations. All audio data in the final corpus have been anonymized.